OpenStudy (anonymous):

Choose the correct simplification of 4^3 / 4^2

OpenStudy (anonymous):



OpenStudy (anonymous):

4

OpenStudy (anonymous):

\[\LARGE {4^3\over 4^2}={4\cdot 4\cdot 4\over 4\cdot 4 }=\] \[\LARGE {\cancel4\cdot \cancel4\cdot 4\over \cancel4\cdot \cancel 4 }=4\] or you just can use the formula: \[\LARGE {x^a\over x^b}=x^{a-b}\] so \[\LARGE {4^3\over 4^2}=4^{3-2}=4^1=4 \]